find the value of n for which the equations (n-1)^2 Uxx - y^2n Uyy= ny^2n-1 is parabola and hyperbola
1
Expert's answer
2020-03-17T16:14:31-0400

"A(x,y)u_{xx}+B(x,y)u_{xy}+C(x,y)u_{yy}=F(x,y,u,u_x,u_y)"


The type of second-order PDE at a point "(x_0,y_0)" depends on the sign of the discriminant defined as


"\Delta(x_o,y_o)=\begin{vmatrix}
 B & 2A \\
 2C & B
\end{vmatrix}"


"\Delta(x_o,y_o)=\begin{vmatrix}
 0 & 2(n-1)^2 \\
 2(-y^{2n})& 0
\end{vmatrix}=4(n-1)^2y^{2n}"

For Parabola;

"4(n-1)^2y^{2n}=0"

"n=1" or "y=0"

For Hyperbola;

"\Delta(x_o,y_o)>0"

"4(n-1)^2y^{2n}>0"

"n\neq1,y\neq0"